# Observational signatures of Rotating compact objects in Plasma space-time

### v1

@@ VII Conclusion
The future low-frequency observations can use our comparison between a wormhole and a black hole to distinguish them from each other however the real environment around such compact objects highly deviates from the current scope of this paper, nevertheless, such studies can help to understand the qualitative nature of the different compact objects. The boundary of the shadow depends on the space-time (considering the geodesics are travelling in the plasma medium in our case) and the gravitational strength of the object Synge (1966); Luminet (1979); Shaikh (2018b). Therefore, in such a scenario, we can understand the qualitative difference between the BH and WH. In our future research endeavors, we plan to delve deeper into the effects of non-uniform plasma densities on both the Kerr black hole and the Teo wormhole. We also plan to compare the results of the parametrised metrics for the black hole by Konoplya et al. Konoplya et al. (2016) with the parameterised wormhole metric by Bronnikov et al.Bronnikov et al. (2021) within the plasma space-time. By examining their behavior under the same plasma profile, we aim to enhance our understanding of their characteristics in more realistic conditions. This investigation will contribute to our comprehension of the intricate interplay between plasma and the propagation of light rays, offering valuable insights into the behavior of these fascinating astrophysical objects.

### v2

@@ VII Conclusion
The future low-frequency observations can use our comparison between a wormhole and a black hole to distinguish them from each other however the real environment around such compact objects highly deviates from the current scope of this paper, nevertheless, such studies can help to understand the qualitative nature of the different compact objects. The boundary of the shadow depends on the space-time (considering the geodesics are travelling in the plasma medium in our case) and the gravitational strength of the object Synge (1966); Luminet (1979); Shaikh (2018b). Therefore, in such a scenario, we can understand the qualitative difference between the BH and WH. In our future research endeavors, we plan to delve deeper into the effects of non-uniform plasma densities on both the Kerr black hole and the Teo wormhole. We also plan to compare the results of the parametrized metrics for the black hole by Konoplya et al. Konoplya et al. (2016) with the parameterized wormhole metric by Bronnikov et al.Bronnikov et al. (2021) within the plasma space-time. By examining their behavior under the same plasma profile, we aim to enhance our understanding of their characteristics in more realistic conditions. This investigation will contribute to our comprehension of the intricate interplay between plasma and the propagation of light rays, offering valuable insights into the behavior of these fascinating astrophysical objects.